Hey I have been having a hard time with the log's here's my three questions,

1. Solve for x; log(5-2x)-log(x-1). Can you also work it out thanks

2. Solve for x; log (x+1)-log(x-1).

3. Log e^x^2=25

Answer by stanbon(48535)

1. Solve for x; log(5-2x)=log(x-1).
5-2x = x-1
3x = 6
x = 2
-------------------------------
2. Solve for x; log(x+1)-log(x-1)=1.
log[(x+1)/(x-1)] = 1
(x+1)/(x-1) = 10
x+1 = 10x-10
8x = 11
x = 11/8
-------------------------------
3. Log e^x^2=125
x^2(log e) = log(125)
x^2 = [log(125)/log e]
x^2 = 4.8283
x = 2.1973...
Comment: I still think this last problem was posted incorrectly.
-------------------------------------
4. log(8)x = 1/3
x = 8^(1/3)
x = 2
================================
Cheers,
Stan H.